Find the integral surface of the equation x2p+y2q+z2=0 passing through z=1,x+y=xy


1
Expert's answer
2021-12-01T13:36:20-0500

"\frac{dx}{x^2}=\frac{dy}{y^2}=\frac{dz}{-z^2}"


"-1\/x=-1\/y+c_1"

"1\/z=-1\/y+c_2"

"F(c_1,c_2)=F(1\/y-1\/x,1\/z+1\/y)=0"


for z=1,x+y=xy :

"c_2=1+1/y\implies y=c"


"c_1=1/y-1/x=\frac{x-y}{xy}=\frac{x-y}{x+y}=\frac{x-c}{x+c}"